# A Proactive Decoy Selection Scheme for Cyber Deception using MITRE ATT&CK

### v1

@@ 1 Introduction
In recent years, the enterprise sector has experienced an increasing number of cyber attacks targeting sensitive data exfiltration, IT infrastructure remote control, data destruction and/or encryption [1]. The involved threat actors leverage a plethora of tactics, techniques and procedures (TTPs) to gain a foothold within the production environment of an organization while remaining undetected, and then persistently achieving a variety of crimes over a long period of time. In this scenario, the adoption of a proactive cyber defense mechanism, such as cyber deception, becomes pivotal in offsetting the strategical advantage of attackers over defenders by weaponizing the latter with decoys that mimic legitimate system components (e.g., text files containing fake login credentials, simulated communication protocols with vulnerabilities, fully emulation of organization services) [2]. The allocation of these software artifacts within the production environment provides defenders the ability to reliably expose stealthy attackers whenever they engage these fake assets and, at the same time, the opportunity to monitor the employed TTPs without affecting the performance of the real system components. The collected information is extremely valuable to build reliable cyber-threat intelligence that can be used to enhance the security posture of the system [3].
@@
Ideally, an effective and practical cyber deception mechanism consists of decoys intentionally exposing vulnerabilities, misconfigurations and/or information that can be exploited by the attacker to progress towards his goal [4]. From a defender standpoint, the selection of decoys with these characteristics is a challenging task as it essentially requires a precise understating of the attacker behavior in order to anticipate his strategy. In this regard, the research activity in the field of cyber deception has partially addressed this problem [5][6][7][8]. In detail, attackers behavior are modelled as random walks in a graph where nodes identify organization assets that can be potential targets and edges indicate the feasible lateral movements of the attacker. Decoys are treated as abstract entities that are placed on non-target nodes to maximize the probability to intercept the attacker path according to some prior knowledge about the adversary attack steps. However, the proposed threat models are not sufficiently accurate to capture the attacker capabilities since they are not supported by empirical observations of the attacker TTPs. Moreover, the considered decoy modeling does not provide any practical insight on what weaknesses these artifacts should expose to reliably lure attackers before they can reach their targets. Consequently, both the applicability and effectiveness of such solutions in real world scenarios is limited.
@@
In light of these observations, we try to bridge the gap towards a more pragmatic cyber deception approach that is driven by an accurate modeling of TTPs employed by real-world attackers in enterprise environments. To accomplish this, we model attackers behavior using enterpriseLang [9], a threat modeling language based on the Meta Attack Language [10]. This domain-specific language provides a formal approach to describe adversaries using the MITRE ATT&CK Enterprise matrix as source of knowledge of real-world TTPs targeting the IT infrastructure of enterprises [11]. This framework provides a record of cyber threats that are extracted from publicly available incident reports in the enterprise sector and that are organized by mapping similar threat actor strategies under the same techniques. The resulting threat model offers a holistic view of the attacker TTPs and related targets and it allows the generation of attack graphs that express the attacker capabilities in terms of techniques and related outcomes on the defender environment. Specifically, each node may correspond to either an ATT&CK technique, whose predecessors indicate the required ATT&CK techniques enabling its execution, or to a technique outcome that reflects the impact of the attacker action on the defender environment.
@@
In the context of cyber deception, we take advantage of this threat model to anticipate the adversarial behavior and proactively select decoys that can be engaged by the attacker using specific ATT&CK techniques. For example, a text file decoy is suitable to lure adversaries employing the ATT&CK technique “Remote System Discovery” which targets the collection of the system elements information from local host files. In this case, a defender that wants to expose an adversary using this technique could locate fake hostnames in a text file located in a monitored system folder that exclusively non-legitimate users would try to access. By extending this reasoning to every ATT&CK techniques, we optimize the selection of decoys by computing the minimum subset of techniques that, if detected by a corresponding number of deceptive artifacts, ensure the interception of all attack paths executed by an attacker to reach a predetermined set of targets. In detail, we resume our contributions as follows:
@@
    We propose a methodology to proactively select decoys in enterprise environments using the MITRE ATT&CK Enterprise matrix. In particular, we leverage the cyber threat information provided by the threat modelling language enterpriseLang to coherently combine ATT&CK techniques according to their execution preconditions. We exploit this data-driven attack path modeling to design a decoy selection scheme tailored to realistic attacker engagement policies.
@@
The remainder of the paper is organized as follows. In Section 2, we discuss the related work. In Section 3, we present some background information. In Section 4, we describe the considered system model, which is composed by the threat and defender models. In Section 5, we formalize the proposed optimization problem that computes the selection of decoys. In Section 6, we present the simulation results. Finally, we draw the conclusion in Section 7.
@@ 2 Related work
Recent work in cyber deception has evolved towards the design of decoys, honeypots and /or honey-tokens solutions to lure TTPs targeting specific defender environments such as enterprise, industrial, internet-of-things. An overview of the proposed solutions and related threat models can be found in [12][13]. Moreover, an increasing number of deception approaches leverage cyber threat intelligence platforms such as MITRE ATT&CK to design and/or evaluate the proposed schemes, ensuring more effective and reliable technologies for practical scenarios [14]. In this context, we highlight our contributions by emphasizing the difference in the methodology that we propose in building a threat model supported by real-world adversary behavior that also captures the execution preconditions between the various techniques.
@@
The authors of [15] and [16] present decoy allocation schemes based on hidden Markov models and partially observable scholastic games, respectively, to predict the most probable attack path of an adversary and place the deceptive artifacts accordingly. The authors of [17] propose a honeypot allocation scheme that leverages core attack graphs to reduce the number of computed attack paths with limited impact on the deception performance, hence improving the deployment scalability. Although these works design practical approximations to render the related algorithms efficient in practical scenarios, the proposed threat models are limited to laterally movements between the various system elements and they ignore the specific TTPs used by attackers to interact and penetrate those components. As a result, they do not provide information about the functionalities that each decoy should implement in order to lure a malicious user. Instead, we leverage a threat modeling language to extend the attacker capabilities beyond abstract lateral movements among system assets using real-world TTPs. This strategy allows us to select decoys tailored to detect specific techniques that implicitly offers practical insights about the deceptive artifact implementation.
@@
The authors of [18] and [19] propose lightweight implementation of deception solutions to detect insider threats within production environments of enterprise systems. Although, their approach is orthogonal to our as we focus on the decoy selection rather than its technical implementation, the lack of a quantitative threat modeling of the attacker behavior prevent a proper performance evaluation, hindering their effectiveness. Differently, we consider a threat model generating attack paths that are constituted by techniques employed in real enterprise environments. This approach enables a proactive decoy allocation, where the defender can design a practical deception strategy by anticipating the adversary behavior.
@@
The authors of [20] and [21] propose several cyber deception strategies based on different methodologies including genetic optimization algorithms to distort attacker perception of IoT environments using decoys. However, they tailor their decoy allocation to reconnaissance and data exfiltration attacks only. Moreover, they do not model the intermediate attack steps to achieve those goals. Consequently, the considered threat model is neither suitable for attack scenarios in the enterprise environments nor it can provide a quantitative characterization of the attacker behavior due to the limited number of considered attack paths.
@@
The authors of [22] propose MTDSim, an framework to evaluate the performance of Moving Target Defense and cyber deception approaches against realistic attackers.The threat model is derived using ATT&CK tactics only, hence it considers the various phases of a cyberattack without providing information about each attack step. Instead, we provide a representation of the attacker behavior by extracting the execution preconditions between techniques from a threat language model using MITRE ATT&CK as TTPs source. This enables the optimization of the amount of selection decoys as we can exploit the execution preconditions between techniques to minimize the number of deceptive artifacts.
@@
In [23] the authors proposes SODA, an orchestration framework that automatically analyzes real-world malware behavior, maps its activities to ATT&CK techniques and design suitable deceptive actions to counteract any malicious software performing those operations on the real system. The authors of [24] provides a performance evaluation of a cyber deception solution for smart grid systems using MITRE ATT&CK. In particular, they map real-world attack paths to ATT&CK techniques and they empirically show that the considered decoy architecture offers a mitigation to each attack step. Differently, the authors of [25] and [26] collect cyber threat intelligence data from deployed honeypots and map the observed attacks to ATT&CK techniques. This information is used to provide a starting point for possible technique mitigations and to cluster similar approaches under a single threat actor, respectively. The aforementioned work employ MITRE ATT&CK to evaluate a-posteriori the effectiveness of the proposed deception strategies in terms of detected techniques. We propose a fundamentally different approach that exploits MITRE ATT&CK to generate attack graphs modeling the various attack steps of an adversary. Based on this, we design the decoy selection to expose the minimum number of techniques used by all feasible attack paths aiming to a predetermined set of targets.
@@ 3 Background
We summarize the main concepts of MITRE ATT&CK framework and EnterpriseLang threat modelling language using the respective original works [11][9]. We remark that these tools are the core elements that we employed to design our decoy selection solution and their description is helpful to provide a better understanding of our contribution in later sections.
@@ 3.1 MITRE ATT&CK framework overview
The ATT&CK Framework provides a structured classification, inspired by the cyber kill chain model [27], of real-world cyber attacks mainly targeting enterprise IT infrastructures. At a high level, ATT&CK can be considered as an adversary behavioral model that represents observed attacks according to the following nomenclature:
@@
    Tactics: they represent the short-term tactical goal of an adversary and essentially correspond to the reason for performing a technique. For example, an adversary want to achieve Credential Access as an intermediate goal in order to accomplish the final target that may be data exfiltration.
@@
    Techniques: they represent how the adversary achieve its tactical goal by describing the general strategy to accomplish that objective. For example, an adversary aiming to achieve Credential Access may search for typical password storage locations within the victim system to obtain the user credential. Each technique may be further split in sub-technique when several existing strategies are available.
Threat actors showing similar behavior are clustered together and are considered as a single adversary, referred to as Group. Their activities are mapped according to the above nomenclature in a matrix where columns are the tactics and rows are the techniques. For each technique, the associated mitigations are also present (when available). As a result, by inspecting a particular group, it is possible to examine how those threat actors operated in the past to achieve specific targets.
@@ 3.2 EnterpriseLang overview
EnterpriseLang is a threat model language that enables the formal description of the adversary behavior in enterprise systems. Its structure is based on the Meta Attack Language (MAL), a general-purpose threat model language that provides syntax and symbol notation to create attacker models for specific domains (e.g., cloud, automotive, internet of things, etc.). For the enterprise scenario, enterpriseLang models the attacker actions using the cyber threat intelligence information provided by the MITRE ATT&CK enterprise matrix. Each one of the 266 ATT&CK technique is translated according to the MAL specification in the following items, which are the main elements constituting the language architecture:
@@
    Attack Step: it represents the action of the adversary and is identified by an ATT&CK technique with the related outcome on the defender environment. Each outcome characterizes the effects of the various ATT&CK techniques on the environment assets (note that some techniques may not have an associated outcome since they do not have an immediate impact on the environment). Examples of technique outcomes includes SSH Credential Interception, Remote Execution, Sensitive Data Collected, Infected Computer.
@@
    Type: it indicates when an adversary is eligible to execute an attack step based on its connections. There are two possible categories, namely AND-type and OR- type. The former identifies an attack step that can be processed by the attacker only if every predecessor attack step is successfully completed. Conversely, the latter identifies an attack step that can be executed as soon as at least one of its predecessor attack step is finalized.
@@ 4 System Model
We consider an organization (i.e., the defender) that employs cyber deception to reliably detects attackers aiming to some targets within its IT enterprise infrastructure. To ensure an effective decoy placement that is tailored to real-world attacker capabilities, it leverages enterpriseLang to model the adversary behavior in terms of ATT&CK techniques and uses such representation to design the decoy selection process. In the next subsections, we formalize the considered threat model and how it is employed by the defender to achieve a practical decoy selection solution. We resume the main notation used in the paper in Table 1.

### v2

@@ 1 Introduction
In recent years, the enterprise sector has experienced an increasing number of cyber attacks targeting sensitive data exfiltration, IT infrastructure remote control, data destruction and/or encryption (Sharif and Mohammed, 2022). The involved threat actors leverage a plethora of tactics, techniques and procedures (TTPs) to gain a foothold within the production environment of an organization while remaining undetected, and then persistently achieving a variety of crimes over a long period of time. In this scenario, the adoption of a proactive cyber defense mechanism, such as cyber deception, becomes pivotal to offset the strategic advantage of attackers over defenders by weaponizing the latter with decoys that mimic legitimate system components (e.g., text files containing fake login credentials, simulated communication protocols with vulnerabilities, fully emulation of organization services) (Wang and Lu, 2018). The allocation of these software artifacts within the production environment provides defenders the ability to reliably expose stealthy attackers whenever they engage these fake assets and, at the same time, the opportunity to monitor the employed TTPs without affecting the performance of the real system components. The information collected is extremely valuable for building reliable cyber threat intelligence that can be used to enhance the security posture of the system (Wagner et al., 2019).
@@
Ideally, an effective and practical cyber deception mechanism consists of decoys intentionally exposing vulnerabilities, misconfigurations and/or information that can be exploited by the attacker to progress towards his goal (Ferguson-Walter et al., 2021). From a defender standpoint, the selection of decoys with these characteristics is a challenging task as it essentially requires a precise understating of the attacker behavior in order to anticipate his strategy. In this regard, the research activity in the field of cyber deception has partially addressed this problem (Sayed et al., 2023; Cromp et al., 2023; Kulkarni et al., 2020; Anwar et al., 2022). In detail, attackers behavior are modeled as random walks in a graph where nodes identify organization assets that can be potential targets and edges indicate the feasible lateral movements of the attacker. Decoys are treated as abstract entities that are placed on non-target nodes to maximize the probability to intercept the attacker path according to some prior knowledge about the adversary attack steps. However, the proposed threat models are not sufficiently accurate to capture the attacker capabilities since they are not supported by empirical observations of the attacker TTPs. Moreover, the considered decoy modeling does not provide any practical insight on what weaknesses these artifacts should expose to reliably lure attackers before they can reach their targets. Consequently, both the applicability and effectiveness of such solutions in real world scenarios is limited.
@@
In light of these observations, we try to bridge the gap towards a more pragmatic cyber deception approach that is driven by an accurate modeling of TTPs employed by real-world attackers in enterprise environments. To accomplish this, we model attackers behavior using enterpriseLang (Xiong et al., 2022), a threat modeling language based on the Meta Attack Language (Johnson et al., 2018). This domain-specific language provides a formal approach to describe adversaries using the MITRE ATT&CK Enterprise matrix as source of knowledge of real-world TTPs targeting the IT infrastructure of enterprises (Strom et al., 2018). This framework provides a record of cyber threats that are extracted from publicly available incident reports in the enterprise sector and that are organized by mapping similar threat actor strategies under the same techniques. The resulting threat model offers a holistic view of the attacker TTPs and related targets and it allows the generation of attack graphs that express the attacker capabilities in terms of techniques and related outcomes on the defender environment. Specifically, each node may correspond to either an ATT&CK technique, whose predecessors indicate the required ATT&CK techniques enabling its execution, or to a technique outcome that reflects the impact of the attacker action on the defender environment.
@@
In the context of cyber deception, we take advantage of this threat model to anticipate adversarial behavior and proactively select decoys that can be engaged by the attacker using specific ATT&CK techniques. For example, a text file decoy is suitable to lure adversaries employing the ATT&CK technique “Remote System Discovery” which targets the collection of the system elements information from local host files. In this case, a defender that wants to expose an adversary using this technique could locate fake hostnames in a text file located in a monitored system folder that exclusively non-legitimate users would try to access. By extending this reasoning to every ATT&CK techniques, we optimize the selection of decoys by computing the minimum subset of techniques that, if detected by a corresponding number of deceptive artifacts, ensure the interception of all attack paths executed by an attacker to reach a predetermined set of targets. In detail, we resume our contributions as follows:
@@
    We propose a methodology to proactively select decoys in enterprise environments using the MITRE ATT&CK Enterprise matrix. In particular, we leverage the cyber threat information provided by the threat modeling language enterpriseLang to coherently combine ATT&CK techniques according to their execution preconditions. We exploit this data-driven attack path modeling to design a decoy selection scheme tailored to realistic attacker engagement policies.
@@
The remainder of the paper is organized as follows. In Section 2, we discuss the related work. In Section 3, we present some background information. In Section 4, we describe the considered system model, which is composed by the threat and defender models. In Section 5, we formalize the proposed optimization problem that computes the selection of decoys. In Section 6, we present the simulation results. Finally, we draw the conclusion in Section Section 7.
@@ 2 Related work
Table 1: Comparison of cyber deception solutions
Reference Focus Threat Model Objective Limitations
Horák et al. (2019); Al Amin et al. (2021); Nguemkam et al. (2023) Decoy allocation Lateral movement Allocation of decoys to maximize the attacker interaction probability Threat model ignores specific TTPs; decoys are not practically characterized
Alohaly et al. (2022); Pagnotta et al. (2023) Decoy implementation Lateral movement Lightweight decoy implementation integrated within production environment to detect insider threats Decoy design is not driven by a quantitative modeling of attacker TTPs
Ge et al. (2021); Rehman et al. (2024) Decoy allocation Reconnaissance, Data exfiltration Allocation of decoys to distort attacker perception of IoT environments Consider only specific attack stages, hence they cannot be extended to more complex attack scenarios composed by multiple attack steps
Brown et al. (2023); Sajid et al. (2021); Mashima (2022); Subhan and Lim (2023); Hobert et al. (2023) Decoy selection ATT&CK techniques Selection of decoys capable to detect specific MITRE ATT&CK techniques Selected decoys ignore possible defense mitigation already in place; techniques preconditions are not exploited to simultaneously intercept multiple attack paths
@@
Recent work in cyber deception has evolved towards the design of decoys, honeypots and /or honey-tokens solutions to lure TTPs targeting specific defender environments such as enterprise, industrial, internet-of-things. An overview of the proposed solutions and related threat models can be found in (Qin et al., 2024; Zhang and Thing, 2021). Moreover, an increasing number of deception approaches leverage cyber threat intelligence platforms such as MITRE ATT&CK to design and/or evaluate the proposed schemes, ensuring more effective and reliable technologies for practical scenarios (Torres et al., 2022). In this context, we highlight our contributions by emphasizing the difference in the methodology that we propose in building a threat model supported by real-world adversary behavior that also captures the execution preconditions between the various techniques. In Table 1, we provide a state-of-the-art summary of the considered works.
@@
The authors of (Horák et al., 2019) and (Al Amin et al., 2021) present decoy allocation schemes based on hidden Markov models and partially observable scholastic games, respectively, to predict the most probable attack path of an adversary and place the deceptive artifacts accordingly. The authors of (Nguemkam et al., 2023) propose a honeypot allocation scheme that leverages core attack graphs to reduce the number of computed attack paths with limited impact on the deception performance, hence improving deployment scalability. Although these works design practical approximations to render the related algorithms efficient in practical scenarios, the proposed threat models are limited to laterally movements between the various system elements and they ignore the specific TTPs used by attackers to interact and penetrate those components. As a result, they do not provide information about the functionalities that each decoy should implement in order to lure a malicious user. Instead, we leverage a threat modeling language to extend the attacker capabilities beyond abstract lateral movements among system assets using real-world TTPs. This strategy allows us to select decoys tailored to detect specific techniques that implicitly offer practical insights about the deceptive artifact implementation.
@@
The authors of (Alohaly et al., 2022) and (Pagnotta et al., 2023) propose a lightweight implementation of deception solutions to detect insider threats within production environments of enterprise systems. Although, their approach is orthogonal to our as we focus on the decoy selection rather than its technical implementation, the lack of a quantitative threat modeling of the attacker behavior prevents a proper performance evaluation, hindering their effectiveness. Differently, we consider a threat model generating attack paths that are constituted by techniques employed in real enterprise environments. This approach enables a proactive decoy allocation, where the defender can design a practical deception strategy anticipating the adversary behavior.
@@
The authors of (Ge et al., 2021) and (Rehman et al., 2024) propose several cyber deception strategies based on different methodologies including genetic optimization algorithms to distort attacker perception of IoT environments using decoys. However, they tailor their decoy allocation to reconnaissance and data exfiltration attacks only. Moreover, they do not model the intermediate attack steps to achieve those goals. Consequently, the considered threat model is neither suitable for attack scenarios in the enterprise environments nor it can provide a quantitative characterization of the attacker behavior due to the limited number of considered attack paths.
@@
The authors of (Brown et al., 2023) propose MTDSim, an framework to evaluate the performance of Moving Target Defense and cyber deception approaches against realistic attackers. The threat model is derived using ATT&CK tactics only, hence it considers the various phases of a cyberattack without providing information about each attack step. Instead, we provide a representation of the attacker behavior by extracting the execution preconditions between techniques from a threat language model using MITRE ATT&CK as TTPs source. This enables the optimization of the amount of selected decoys as we can exploit the execution preconditions between techniques to minimize the number of deceptive artifacts.
@@
In (Sajid et al., 2021), the authors propose SODA, an orchestration framework that automatically analyzes real-world malware behavior, maps its activities to ATT&CK techniques and design suitable deceptive actions to counteract any malicious software performing those operations on the real system. The authors of (Mashima, 2022) provide a performance evaluation of a cyber deception solution for smart grid systems using MITRE ATT&CK. In particular, they map real-world attack paths to ATT&CK techniques and they empirically show that the considered decoy architecture offers a mitigation to each attack step. Differently, the authors of (Subhan and Lim, 2023) and (Hobert et al., 2023) collect cyber threat intelligence data from deployed honeypots and map the observed attacks to ATT&CK techniques. This information is used to provide a starting point for possible technique mitigations and to cluster similar approaches under a single threat actor, respectively. The aforementioned work employ MITRE ATT&CK to evaluate a-posteriori the effectiveness of the proposed deception strategies in terms of detected techniques. We propose a fundamentally different approach that exploits MITRE ATT&CK to generate attack graphs modeling the various attack steps of an adversary. Based on this, we design the decoy selection to expose the minimum number of techniques used by all feasible attack paths aiming to a predetermined set of targets.
@@ 3 Background
We summarize the main concepts of MITRE ATT&CK framework and EnterpriseLang threat modeling language using the respective original works (Strom et al., 2018; Xiong et al., 2022). We remark that these tools are the core elements that we employed to design our decoy selection solution and their description is helpful to provide a better understanding of our contribution in later sections.
@@ 3.1 MITRE ATT&CK framework overview
The MITRE ATT&CK Framework provides a structured classification, inspired by the cyber kill chain model (Hutchins et al., 2011), of real-world cyber attacks mainly targeting enterprise IT infrastructures. At a high level, ATT&CK can be considered as an adversary behavioral model that represents observed attacks according to the following nomenclature:
@@
    Tactics: they represent the short-term tactical goal of an adversary and essentially correspond to the reason for performing a technique. For example, an adversary wants to achieve Credential Access as an intermediate goal in order to accomplish the final target that may be data exfiltration.
@@
    Techniques: they represent how the adversary achieves its tactical goal by describing the general strategy to accomplish that objective. For example, an adversary aiming to achieve Credential Access may search for typical password storage locations within the victim system to obtain the user credential. Each technique may be further split in sub-techniques when several existing strategies are available.
Threat actors showing similar behavior are clustered and considered as a single adversary, referred to as Group. Their activities are mapped according to the above nomenclature in a matrix where columns are the tactics and rows are the techniques. For each technique, the associated mitigations are also present (when available). As a result, by inspecting a particular group, it is possible to examine how those threat actors operated in the past to achieve specific targets.
@@ 3.2 EnterpriseLang overview
EnterpriseLang is a threat modeling language that enables the formal description of the adversary behavior in enterprise systems. Its structure is based on the Meta Attack Language (MAL), a general-purpose threat model language that provides syntax and symbol notation to create attacker models for specific domains (e.g., cloud, automotive, internet of things, etc.). For the enterprise scenario, enterpriseLang models the attacker actions using the cyber threat intelligence information provided by the MITRE ATT&CK enterprise matrix. Each one of the 266 ATT&CK technique is translated according to the MAL specification in the following items, which are the main elements constituting the language architecture:
@@
    Attack Step: it represents the action of the adversary and is identified by an ATT&CK technique with the related outcome on the defender environment. Each outcome characterizes the effects of the various ATT&CK techniques on the environment assets (note that some techniques may not have an associated outcome since they do not have an immediate impact on the environment). Examples of technique outcomes include SSH Credential Interception, Remote Execution, Sensitive Data Collected, Infected Computer.
@@
    Type: it indicates when an adversary is eligible to execute an attack step based on its connections. There are two possible type categories, namely AND type and OR type. The former identifies an attack step that can be processed by the attacker only if every predecessor attack step is successfully completed. Conversely, the latter identifies an attack step that can be executed as soon as at least one of its predecessor attack step is finalized.
@@ 4 System Model
Figure 1: Visual representation of the proposed threat model. The attacker capabilities are expressed as a subset of ATT&CK techniques (red squares) of the MITRE ATT&CK Enterprise matrix, where techniques within the same column belong to the same tactic. The considered techniques are combined according to their execution preconditions using Connection and Type information encoded within enterpriseLang. Similarly, the attack path representation is further augmented with the technique outcomes and possible defense mitigations using Attack Step and Defense mitigations information, respectively.
We consider an organization (i.e., the defender) that employs cyber deception to reliably detect attackers aiming to some targets within his IT infrastructure. To ensure an effective decoy placement that is tailored to real-world attacker capabilities, it leverages enterpriseLang to model the adversary behavior in terms of ATT&CK techniques and uses such representation to design the decoy selection process. In the next subsections, we formalize the threat model considered and how it is used by the defender to achieve a practical decoy selection solution. We resume the main notation used in the paper in Table 2.